# InvertAvatar: Incremental GAN Inversion for Generalized Head Avatars

### v2

@@ Abstract
While high fidelity and efficiency are central to the creation of digital head avatars, recent methods relying on 2D or 3D generative models often experience limitations such as shape distortion, expression inaccuracy, and identity flickering. Additionally, existing one-shot inversion techniques fail to fully leverage multiple input images for detailed feature extraction. We propose a novel framework, Incremental 3D GAN Inversion, that enhances avatar reconstruction performance using an algorithm designed to increase the fidelity from multiple frames, resulting in improved reconstruction quality proportional to frame count. Our method introduces a unique animatable 3D GAN prior with two crucial modifications for enhanced expression controllability alongside an innovative neural texture encoder that categorizes texture feature spaces based on UV parameterization. Differentiating from traditional techniques, our architecture emphasizes pixel-aligned image-to-image translation, mitigating the need to learn correspondences between observation and canonical spaces. Furthermore, we incorporate ConvGRU-based recurrent networks for temporal data aggregation from multiple frames, boosting geometry and texture detail reconstruction. The proposed paradigm demonstrates state-of-the-art performance on one-shot and few-shot avatar animation tasks.
@@
Figure 1: Given single or more source images, our method rapidly reconstructs photorealistic 3D facial avatars under one second, enabling precise control over full-head rotations and subtle expressions like lopsided grins and eye gazes.
footnotetext: *Equal Contribution
1 Introduction
@@
The creation of digital head avatars [16, 59, 19, 57, 60, 3, 11, 13, 35, 48] has long been a task of interest in the fields of computer vision and graphics. The ability to interactively generate photo-realistic head avatars opens up new possibilities in areas such as augmented/virtual reality (AR/VR), 3D telepresence, and video conferencing. The primary aims when developing a digital head avatar are two-fold: high fidelity and efficiency. High fidelity involves preserving intricate details from input images, reconstructing occluded facial regions (especially in profile views), and accurately modeling dynamic features such as facial expressions, all while maintaining consistent identity throughout animation. Efficiency demands a model that can adapt to varied unseen identities and motions without additional optimization at the inference stage.
@@
Recent literature showcases numerous attempts to create avatars from a single input image and perform animation with video frames. Several 2D generative models [11, 13, 35, 48] perform image animation by incorporating the 3D Morphable Face Models (3DMM) [7] into the portrait synthesis. These 2D-based methods achieve photo-realism but suffer from shape distortion during large motion due to a lack of geometry constraints.
@@
Towards better view consistency, many recent efforts [4, 31] leverage 3D GAN prior and decouple motion and appearance in latent space with video data. The main drawback of this paradigm is that the motion and appearance are naturally entangled in 3D GAN which leads to expression inaccuracy and identity flickering during animation. The other line of works [41, 40, 29, 51] directly learn a controllable 3D GAN from only 2D image collections, and perform GAN inversion to restore an avatar from one single image. The methods employing the controllable generative prior demonstrate efficiency, yet they are constrained by their dependency on a singular source image. A single image often falls short in fully representing the subject owing to factors such as occlusions and limited pose information. Utilizing diverse source images enriches appearance data, thereby diminishing the extent of hallucination required by an image generator. However, these one-shot inversion techniques do not directly enable the aggregation of partial observations across distinct frames to reconstruct more personalized details. Consequently, this represents a significant limitation to their capacity for detailed feature extraction.
@@
To accumulate temporal data from numerous frames, we further modify the one-shot inversion architecture to accommodate multiple inputs by leveraging ConvGRU-based [5] recurrent networks. This unique recurrent mechanism allows a flexible number of inputs and autonomously determines which information should be retained or discarded within streaming frames, thereby bolstering geometry and texture detail reconstruction capabilities.
@@
  • We provide a new paradigm for efficient avatar reconstruction for streaming data based on a concept called incremental 3D GAN inversion. The paradigm supports one to multiple source images inputs and achieves state-of-the-art performance on one-shot and few-shot facial avatar animation tasks.
@@
2 Related Works
2.1 3D GAN inversion
@@ 2 Related Works
Building upon the achievements of 2D Generative Adversarial Network (GAN) inversion for image editing and manipulation [36, 45, 12, 2, 49], current 3D GAN inversion methodologies [26, 39, 30] perform a projection of specific images onto multiple instances within the pre-trained StyleGAN2 latent space [23, 1]. While the global latent space enhances 3D-aware portrait editing abilities, it results in a trade-off between reconstruction accuracy and editability, thereby complicating the accurate reconstruction of input images. Consequently, contemporary 3D GAN inversion techniques necessitate an estimated camera pose and slight generator weight adjustment [37, 14, 52] during testing to reconstruct out-of-domain input images accurately. In contrast to latent code inversion or tedious generator fine-tuning, recent efforts have shifted towards directly applying 3D GAN inversion on a triplane architecture for 3D GANs. For instance, TriPlaneNet [6] presents a triplane encoder while Live3DPortrait [46] integrates a Vision Transformer architecture into the triplane encoder, both showing substantial enhancements in robustness and 3D inversion quality for out-of-domain images. Furthermore, Control3Diff [20] suggests a diffusion-based tri-plane encoder that demonstrates multimodal control capabilities.
2.2 2D Facial animation with 3DMM
@@
Blanz et al. [7] pioneered the modeling of facial texture and shape in vector spaces, leading to the development of the 3D Morphable Model (3DMM). Subsequent extensions of the 3DMM concept include full-head PCA models [9, 34] and blend-shape models [28], which have been extensively studied and widely implemented in facial animation tasks [18, 43]. While these methods leverage the benefits of 3DMM to accurately and continuously model facial part deformations, they grapple with representing non-facial areas that are beyond the scope of 3DMM, such as hair, teeth, eyes, and body. These methods also tend to lack fine-grained facial details. To compensate for these limitations and generate more realistic facial details, subsequent research [25, 54, 44, 17, 15, 42, 11, 13, 35, 48] has applied learned techniques atop 3DMM renderings. For instance, DiscoFaceGAN [11] maps 3DMM parameters into a latent space, subsequently decoupling them through imitative-contrastive learning. While these 2D-based approaches have successfully facilitated efficient and photorealistic animation, they do not model 3D geometry and hence fail to maintain stringent 3D consistency during significant head pose alterations.
2.3 Neural 3D head avatars
@@
For strict 3D consistency, recent efforts [16, 59, 19, 57, 3] incorporate 3DMM into volumetric representations [32, 38, 21, 33] to achieve view-consistent facial animation. Although these methods demonstrate realistic reconstruction results, they inefficiently learn separate networks for different identities and require thousands of frames from a specific individual for training. Another line of works [40, 31, 56, 29, 4] focus on reconstructing 3D portraits in one-shot or few-shot manner incorporating 3D GAN prior or training in a generative manner. Next3D [40] learns a controllable 3D head avatar from random noise and reconstructs an avatar by PTI inversion. Li et al. [29] proposes to encode input images into a canonical branch for coarse geometry and texture, and introduces an additional appearance branch that captures personal details and an expression branch that modifies the reconstruction to the desired expression. These methods show efficient promising one-shot reconstruction but none of them are architected to accommodate the utilization of multiple source images as input. Contrastingly, our proposed framework is designed to leverage the complementary information inherent in multiple input images, thereby facilitating superior reconstruction quality.
@@
Figure 2: Our avatar reconstruction pipeline, applicable to a single or multi-image sequence, comprises two stages. The coarse stage inverts the first frame in the GAN prior’s $W+$ latent space with $E_{latent}$, forming an initial avatar. The fine stage employs recurrent networks $E_{tex\_rec}$ and $E_{tri\_rec}$ to aggregate temporal information, incrementally refining a high-fidelity avatar in two canonical feature spaces. The architecture of our advanced animatable 3D generative model is depicted in the right box.
3 Method
@@
Our method innovatively harnesses the prior knowledge embedded in an animatable 3D generative model, integrating it with an incremental inversion architecture to enable rapid and high-fidelity 3D avatar reconstruction. Utilizing either a single source image or a short sequence of dynamic frames from a source video, our system is capable of reconstructing a 3D facial avatar, which is readily animatable through a 3D Morphable Face Model (3DMM) [47]. In Sec. 3.1, we introduce our compact yet highly expressive animatable GAN, which provides a strong 3D prior. Sec. 3.2 details our one-shot learning-based avatar reconstruction method, intricately designed to capitalize on the aforementioned GAN prior. Conclusively, in Section 3.3, we describe the employment of RNN-based temporal aggregation networks, which are instrumental in enhancing the shape and appearance precision of the reconstructed avatar by harnessing temporal data.
3.1 Compact and Expressive Animatable 3D GAN
Drawing from Next3D [40] for our 3D facial avatar generator, we identified its drawbacks that restrict wider usability. Its complexity and numerous network parameters complicate inversion-friendly feature space creation, lead to high memory overhead, limit neural rendering resolution, and introduce view inconsistency. It also struggles with accurate animation according to the FLAME model. To address these, we proposed two key architectural changes enhancing animation precision and view consistency, reducing model size by half while retaining synthesis quality.
Compact face synthesis module. Next3D transforms neural texture into orthogonal views and forms tri-plane features. To compensate for the FLAME model’s missing inner mouth area, Next3D utilizes a UNet to supplement and reintegrate this texture, using another UNet for post-blending. These operations are redundant and lead to heavy parameter size, so we alternately propose a more compact architecture, maintaining the insight of completing inner mouth feature conditioning on global information, but removing heavy networks for specific mouth synthesis. Specifically, we propose a face synthesis module $G_{face}$, which is conditioned on the multi-scale rasterized neural textures. The multi-scale neural textures are rasterized from texture space into tri-plane space and then integrated with $G_{face}$’s corresponding feature maps through alpha blending. Conditioned on such global information, $G_{face}$ is precisely controlled by the deformable mesh model, and gradually inpaint inner mouth features in the forward pass layer by layer. Additionally, our more memory-efficient and compact model design allows for neural rendering resolution of $128^{2}$, significantly enhancing view-consistency.
@@ 3.1 Compact and Expressive Animatable 3D GAN
More expressive facial parametric model. The expressiveness of the facial parametric model is paramount for aligning the conditional morphable model with synthesized images. A misalignment between 3DMM and 2D image expressions can weaken their correlation during adversarial training, impeding the correct deformation of generated images. To rectify this, we employ Faceverse [47], a more expressive parametric model than the FLAME model [27]. With its extensive shape and expression bases, Faceverse accurately fits subtle expression changes in 2D images. During training, supplying the discriminator with the tracked landmarks map bolsters the consistency between synthesized images and the conditional mesh.
@@
3.2 One-shot Avatar Reconstruction
To extract maximum personal information from a single image with the reconstructed avatar, we introduce an effective one-shot 3D GAN inversion framework. This framework utilizes a two-stage coarse-to-fine inversion architecture that carries out inversions in the latent and feature spaces of our animatable generative model separately. Due to the disentangled nature of the dynamic face and static head features, we divide the feature space into two specific canonical spaces: $F_{tex}$ defined by $G_{tex}$ (the canonical texture space), and $F_{tri}$ defined by $G_{static}$ (the canonical tri-plane space). The ultimate goal then becomes learning the mapping from the input portrait image to the facial avatar represented in these latent and canonical feature spaces.
In the coarse stage, we introduce a latent encoder $E_{latent}$ to project the image to the $W+$ latent space of the pretrained GAN. Specifically, given a source image $I$, we embed an inverted latent code $\hat{w}=E_{latent}(I)$ and initially generate a coarse facial avatar, represented by canonical features $F_{tex_{w}}=G_{tex}(\hat{w})$ and $F_{tri_{w}}=G_{static}(\hat{w})$. With the estimated pose and expression of the source image $I$, we can synthesize the fake image $\hat{I}=R(G_{face}(F_{tex_{w}}),F_{tri_{w}})$, where $R$ is the rendering block.
In the second stage, with the aim to offset the information loss instigated by $E_{latent}$, we formulate two separate image encoders $E_{tex}$ and $E_{tri}$. These are specifically designed to refine the feature maps in each of the two canonical spaces independently. Different from existing methods that rely on networks to map a posed image to the canonical tri-plane representation, benefiting from the neural texture architecture of our GAN prior, the face appearance represented by texture feature space is explicitly defined on the UV parameterization. Hence we design a U-Net network $E_{tex}$ to translate the observation to the canonical texture feature space in the UV domain, as shown in Fig. 2. Such a pixel-aligned image-to-image translation architecture helps eliminate the learning burden of correspondences between observation space and canonical space, making it concentrate on recovering fine details, as demonstrated in Sec. 4.3.2. Specifically, the input of the $E_{tex}$ contains three maps, the source image $I$ and the residual map $\Delta{I}=I-\hat{I}$ that are projected to the UV plane, and a mask indicating the visibility region of the face. $E_{tex}$ outputs multi-resolution feature offsets $F_{tex_{off}}$ to compensate for the texture features $F_{tex_{w}}$ obtained in the first stage.
As for the static tri-plane features, we introduce an image-to-plane encoder $E_{tri}$ to predict multi-scale feature maps. Specifically, the multi-resolution features $F_{tri-SFT}$ output by $E_{tri}$ are used to spatially modulate $G_{static}$ with the CS-SFT layer [50] in a coarse-to-fine manner.
The training for all image encoders can also be divided into two stages. Firstly, we only train $E_{latent}$, with the loss terms following [45, 36]. Secondly, we simultaneously train $E_{tex}$ and $E_{tri}$. Referring to Live 3D Portrait [46], our training set includes the synthetic data and we supervise the training with many intermediate feature maps, as well as an adversarial loss to generate photo-realistic results. Please refer to our supplemental material for more details.
3.3 Incremental GAN inversion

### v3

@@ Abstract.
While high fidelity and efficiency are central to the creation of digital head avatars, recent methods relying on 2D or 3D generative models often experience limitations such as shape distortion, expression inaccuracy, and identity flickering. Additionally, existing one-shot inversion techniques fail to fully leverage multiple input images for detailed feature extraction. We propose a novel framework, Incremental 3D GAN Inversion, that enhances avatar reconstruction performance using an algorithm designed to increase the fidelity from multiple frames, resulting in improved reconstruction quality proportional to frame count. Our method introduces a unique animatable 3D GAN prior with two crucial modifications for enhanced expression controllability alongside an innovative neural texture encoder that categorizes texture feature spaces based on UV parameterization. Differentiating from traditional techniques, our architecture emphasizes pixel-aligned image-to-image translation, mitigating the need to learn correspondences between observation and canonical spaces. Furthermore, we incorporate ConvGRU-based recurrent networks for temporal data aggregation from multiple frames, boosting geometry and texture detail reconstruction. The proposed paradigm demonstrates state-of-the-art performance on one-shot and few-shot avatar animation tasks. Code will be available at https://github.com/XChenZ/invertAvatar.
3D head avatar, one-shot reconstruction, few-shot reconstruction, recurrent neural network, GAN inversion
submissionid: 859journalyear: 2024copyright: rightsretainedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657478isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Motion processingccs: Computing methodologies Image-based rendering
@@
Figure 1. Given single or more source images, our method rapidly reconstructs photorealistic 3D facial avatars under one second, enabling precise control over full-head rotations and subtle expressions like lopsided grins and eye gazes.
1. Introduction
@@
The creation of digital head avatars (Gafni et al., 2021; Zheng et al., 2022; Grassal et al., 2022; Zhang et al., 2022; Zheng et al., 2023; Athar et al., 2022; Deng et al., 2020; Doukas et al., 2021; Ren et al., 2021; Wang et al., 2023) has long been a task of interest in the fields of computer vision and graphics. The ability to interactively generate photo-realistic head avatars opens up new possibilities in areas such as augmented/virtual reality (AR/VR), 3D telepresence, and video conferencing. The primary aims when developing a digital head avatar are two-fold: high fidelity and efficiency. High fidelity involves preserving intricate details from input images, reconstructing occluded facial regions (especially in profile views), and accurately modeling dynamic features such as facial expressions, while maintaining consistent identity throughout animation. Efficiency demands a model that can adapt to varied unseen identities and motions without additional optimization at the inference stage.
@@
Recent literature showcases numerous attempts to create avatars from a single input image and perform animation with video frames. Several 2D generative models (Deng et al., 2020; Doukas et al., 2021; Ren et al., 2021; Wang et al., 2023) perform image animation by incorporating the 3D Morphable Face Models (3DMM) (Blanz and Vetter, 1999) into the portrait synthesis. These 2D-based methods achieve photo-realism but suffer from shape distortion during large motion due to a lack of geometry constraints.
@@
Towards better view consistency, many recent efforts (Bai et al., 2023; Ma et al., 2023) leverage 3D GAN prior and decouple motion and appearance in latent space with video data. The main drawback of this paradigm is that the motion and appearance are naturally entangled in 3D GAN which leads to expression inaccuracy and identity flickering during animation. The other line of works (Tang et al., 2023; Sun et al., 2023; Li et al., 2023a; Wu et al., 2022) directly learn a controllable 3D GAN from only 2D image collections, and perform GAN inversion to restore an avatar from one single image. The methods employing the controllable generative prior demonstrate efficiency, yet they are constrained by their dependency on a singular source image. A single image often falls short in fully representing the subject owing to factors such as occlusions and limited pose information. Utilizing diverse source images enriches appearance data, thereby diminishing the extent of hallucination required by an image generator. However, these one-shot inversion techniques do not directly enable the aggregation of partial observations across distinct frames to reconstruct more personalized details. Consequently, this represents a significant limitation to their capacity for detailed feature extraction.
@@
To accumulate temporal data from numerous frames, we further modify the one-shot inversion architecture to accommodate multiple inputs by leveraging ConvGRU-based (Ballas et al., 2015) recurrent networks. This unique recurrent mechanism allows a flexible number of inputs and autonomously determines which information should be retained or discarded within streaming frames, thereby bolstering geometry and texture detail reconstruction capabilities.
@@
  • We provide a new paradigm for efficient avatar reconstruction for streaming data based on a concept called incremental 3D GAN inversion. The paradigm supports one to multiple source images inputs and achieves state-of-the-art performance on one- and few-shot facial avatar animation tasks.
@@
2. Related Works
2.1. 3D GAN inversion
@@ 2. Related Works
Building upon the achievements of 2D Generative Adversarial Network (GAN) inversion for image editing and manipulation (Richardson et al., 2021; Tov et al., 2021; Dinh et al., 2022; Alaluf et al., 2021; Wang et al., 2022b), current 3D GAN inversion methodologies (Ko et al., 2023; Sun et al., 2022; Lin et al., 2022) perform a projection of specific images onto multiple instances within the pre-trained StyleGAN2 latent space (Karras et al., 2020; Abdal et al., 2019). While the global latent space enhances 3D-aware portrait editing abilities, it results in a trade-off between reconstruction accuracy and editability, thereby complicating the accurate reconstruction of input images. Consequently, contemporary 3D GAN inversion techniques necessitate an estimated camera pose and slight generator weight adjustment (Roich et al., 2021; Feng et al., 2022; Xie et al., 2023) during testing to reconstruct out-of-domain input images accurately. In contrast to latent code inversion or tedious generator fine-tuning, recent efforts have shifted towards directly applying 3D GAN inversion on a triplane architecture for 3D GANs. For instance, TriPlaneNet (Bhattarai et al., 2023) presents a triplane encoder while Live3DPortrait (Trevithick et al., 2023) integrates a Vision Transformer architecture into the triplane encoder, both showing substantial enhancements in robustness and 3D inversion quality for out-of-domain images. Furthermore, Control3Diff (Gu et al., 2023) suggests a diffusion-based tri-plane encoder that demonstrates multimodal control capabilities.
2.2. 2D Facial animation
@@
Leveraging the benefits of 3DMM, 2D facial animation tasks (Gecer et al., 2019; Thies et al., 2016) are able to accurately and continuously model facial part deformations. However, these methods tend to lack fine-grained facial details, and grapple with representing non-facial areas that are beyond the scope of 3DMM, such as hair, teeth, eyes, and body. To compensate for these limitations and generate more realistic facial details, subsequent research (Kim et al., 2018; Xu et al., 2020; Thies et al., 2019; Gecer et al., 2018; Fried et al., 2019; Tewari et al., 2020; Deng et al., 2020; Doukas et al., 2021; Ren et al., 2021; Wang et al., 2023) has applied learned techniques atop 3DMM renderings. Getting rid of 3DMM, some methods (Burkov et al., 2020; Liang et al., 2022; Drobyshev et al., 2022; Wang et al., 2021b) are trained on the large-scale face video datasets containing rich identities and expressions, thus can be generalized to unseen motion and identity given just a single input image. While these 2D-based approaches have successfully facilitated efficient and photorealistic animation, they do not model 3D geometry and hence fail to maintain stringent 3D consistency during significant head pose alterations.
2.3. Neural 3D head avatars
@@
For strict 3D consistency, recent efforts (Gafni et al., 2021; Zheng et al., 2022; Grassal et al., 2022; Zhang et al., 2022; Athar et al., 2022; Xu et al., 2023b; Zhao et al., 2023; Chen et al., 2024; Xu et al., 2024) incorporate 3DMM into implicit representations (Mildenhall et al., 2020; Sitzmann et al., 2019; Kanazawa et al., 2018; Peng et al., 2020; Kerbl et al., 2023) to achieve view-consistent facial animation. Although these methods demonstrate realistic reconstruction results, they inefficiently learn separate networks for different identities and require thousands of frames from a specific individual for training. Another line of works (Sun et al., 2023; Ma et al., 2023; Yu et al., 2023; Li et al., 2023a; Bai et al., 2023; Li et al., 2023b; Xu et al., 2023a) focus on reconstructing 3D portraits in one-shot or few-shot manner incorporating 3D GAN prior or training in a generative manner. Next3D (Sun et al., 2023) learns a controllable 3D head avatar from random noise and reconstructs an avatar by PTI inversion.  (Li et al., 2023a) proposes to encode input images into a canonical branch for coarse geometry and texture, and introduces an additional appearance branch that captures personal details and an expression branch that modifies the reconstruction to the desired expression. These methods show efficient promising one-shot reconstruction but none of them are architected to accommodate the utilization of multiple source images as input. Contrastingly, our proposed framework is designed to leverage the complementary information inherent in multiple input images, thereby facilitating superior reconstruction quality.
Figure 2. The architecture review of Next3D, along with the visualization of the modifications in our adopted 3D generative model.
@@
Figure 3. The left part illustrates our two-stage avatar reconstruction pipeline. The coarse stage ”Latent Space Inversion” (Sec3.3) inverts the first frame in the GAN prior’s $W+$ latent space with $E_{latent}$, forming an initial avatar. The fine stage performs offset prediction in canonical feature spaces (Sec3.3) and we specifically design recurrent networks $E_{tex\_rec}$ and $E_{tri\_rec}$ to aggregate temporal information, incrementally refining a high-fidelity avatar (Sec3.4). The architecture of our advanced animatable 3D generative model is depicted in the right box.
3. Method
@@
Our method innovatively harnesses the prior knowledge embedded in an animatable 3D generative model, integrating it with an incremental inversion architecture to enable rapid and high-fidelity 3D avatar reconstruction. Utilizing either a single source image or a short sequence of dynamic frames from a source video, our system is capable of reconstructing a 3D facial avatar, which is readily animatable through a 3D Morphable Face Model (3DMM) (Wang et al., 2022a). In Sec. 3.2, we introduce our compact yet highly expressive animatable GAN, which provides a strong 3D prior. Sec. 3.3 details our one-shot learning-based avatar reconstruction method, intricately designed to capitalize on the aforementioned GAN prior. Conclusively, in Section 3.4, we describe the employment of RNN-based temporal aggregation networks, which are instrumental in enhancing the shape and appearance precision of the reconstructed avatar by harnessing temporal data.
3.1. Preliminary: Next3D
Our chosen framework for animatable 3D GAN is Next3D (Sun et al., 2023), which currently stands as the SOTA in generating high-quality, 3D consistent facial avatars from unstructured 2D images. The innovation within Next3D lies in its proposed 3D representation, termed Generative Texture-Rasterized Tri-planes. Illustrated in Fig. 2, given a randomly sampled latent code $z$, the mapping network maps $z$ to an intermediate latent code $w$. This intermediate latent code modulates a StyleGAN generator to synthesize Generative Neural Textures atop parametric mesh templates. Subsequently, these are projected onto three orthogonally oriented feature planes via rasterization of a conditional facial 3DMM, thus establishing a tri-plane feature schema for volume rendering and 2D super-resolution ($G_{superres}$). This dual approach combines the nuanced expression control afforded by mesh-guided explicit deformation with the adaptive prowess of implicit volumetric representation. For more details please refer to (Sun et al., 2023).
3.2. Compact and Expressive Animatable 3D GAN
Drawing from Next3D for our 3D facial avatar generator, we make the following improvements for our avatar reconstruction task. First, since 3DMM does not contain the inner mouth area, Next3D supplements and reintegrates the inner mouth features by two additional UNets, which leads to additional complexity and meory burden. To improve this, as shown in Fig. 2, we replace the two UNets with a single generator $G_{face}$ by conditioning $G_{face}$ with the multi-scale rasterized neural texture feature maps, which allows the inner mouth features to be gradually generated through multiple convolutional layers. Consequently, our more memory-efficient model design allows for larger neural rendering resolution of $128^{2}$.
@@ 3.2. Compact and Expressive Animatable 3D GAN
Besides, we practically notice that the misalignment between conditional 3DMM and 2D image expressions can weaken their correlation during adversarial training, impeding the correct deformation of generated images. To rectify this, we employ Faceverse (Wang et al., 2022a), a parametric model with more expression bases, to accurately fit subtle expression changes in 2D images. During training, we also supply the discriminator with the tracked landmarks map to bolster the consistency between synthesized images and the conditional mesh.
@@
3.3. One-shot Avatar Reconstruction
To extract maximum personal information from a single image with the reconstructed avatar, we introduce an effective one-shot 3D GAN inversion framework. This framework utilizes a coarse-to-fine inversion architecture that carries out inversions in the latent and feature spaces of our animatable generative model separately. Due to the disentangled nature of the dynamic face and static head features, we divide the feature space into two specific canonical spaces: $F_{tex}$ defined by $G_{tex}$ (the canonical texture space), and $F_{tri}$ defined by $G_{static}$ (the canonical tri-plane space). The ultimate goal then becomes learning the mapping from the input portrait image to the facial avatar represented in these latent and canonical feature spaces.
In the first stage, we introduce a latent encoder $E_{latent}$ to project the image to the $W+$ latent space of the pretrained GAN. Specifically, given a source image $I$, we embed an inverted latent code $\hat{w}=E_{latent}(I)$ and initially generate a coarse facial avatar, represented by canonical features $F_{tex_{w}}=G_{tex}(\hat{w})$ and $F_{tri_{w}}=G_{static}(\hat{w})$. With the estimated pose and expression of the source image $I$, we can synthesize the fake image $\hat{I}=R(G_{face}(F_{tex_{w}}),F_{tri_{w}})$, where $R$ is the rendering block.
[…]
As for the static tri-plane features, we introduce an image-to-plane encoder $E_{tri}$ to predict multi-scale feature maps. Specifically, the multi-resolution features $F_{tri-SFT}$ output by $E_{tri}$ are used to spatially modulate $G_{static}$ with the Channel-Split SFT [^2] Channel-Split Spatial Feature Transform is a layer tailored for the StyleGAN2 backbone to perform spatial modulation while preserving the generative prior. layer (Wang et al., 2021a) in a coarse-to-fine manner.
The training for all image encoders can also be divided into two stages. Firstly, we only train $E_{latent}$, with the loss terms following (Tov et al., 2021; Richardson et al., 2021). Secondly, we simultaneously train $E_{tex}$ and $E_{tri}$. Referring to Live 3D Portrait (Trevithick et al., 2023), our training set includes the synthetic data and we supervise the training with many intermediate feature maps, as well as an adversarial loss to generate photo-realistic results. Please refer to our supplemental material for more details.
3.4. Incremental GAN inversion